# Observational Features of Thin Accretion Disk Around Rotating Regular Black Hole

### v1

@@ IV.1 Ray-tracing method
In order to obtain the black hole image, we will apply the ray-tracing method. By shooting light rays from an observer image plane backward to the vicinity of the black hole, we could compute the resulting intensity in the image plane based on how the light rays intersect with the emission source (or the accretion disk) around the black holes.
@@
$\alpha=-\frac{\lambda}{\sin\theta_{o}},\quad\beta=\pm_{o}\sqrt{\eta+a^{2}\cos^{2}\theta_{o}-\lambda^{2}\cot^{2}\theta_{o}}$ (24)
@@
$\lambda=-\alpha\sin\theta_{o},\quad\eta=\left(\alpha^{2}-a^{2}\right)\cos^{2}\theta_{0}+\beta^{2}$ (25)
The conserved quantities corresponding to photon shells or bounded orbits $(\tilde{\lambda},\tilde{\eta})$ (see Eq.17-19) will result in the critical curve $\mathcal{C}=\{\tilde{\alpha},\tilde{\beta}\}$ in the image plane since they are not stable. The enclose region in the image plane by this critical curve is usually denoted as the Black Hole Shadow. This critical curve will determine the behavior of light rays when we shoot backward from the image plane to the rotating regular black hole. For the light rays inside the critical curve, they will fall into black hole. While the light rays outside the critical curve are all scattered back to infinity, which means that they have a turning point in their trajectory. The radial turning point is the largest root of the radial potential $r_{4}$, and it is greater than the event horizon.
The relation Eq.(25) is not enough to determine the source coordinates of a photon. For a given black hole spin $a$, observer radial location $r_{o}$ and inclination angle $\theta_{o}$, our goal is to know the relation between the coordinates of the image plane $(\alpha,\beta)$ and the source coordinates $(r_{s},\theta_{s},\phi_{s},t_{s})$ where this photon emitted from. To do so, we need parameterize the photon trajectories by integrating the geodesic equations (3-6). The differential geodesic equation could recast into integral form,
@@
$\Delta\phi:=\phi_{o}-\phi_{s}$ $=I_{\phi}+\lambda G_{\phi},$ (27)
$\Delta t:=t_{o}-t_{s}$ $=I_{t}+a^{2}G_{t},$ (28)
@@
$I_{r}$ $=f_{r_{s}}^{r_{o}}\frac{\mathrm{~{}d}r}{\pm_{r}\sqrt{\mathcal{R}(r)}}$ (29)
$I_{\phi}$ $=f_{r_{s}}^{r_{o}}\frac{a(2Mre^{-k/r}-a\lambda)}{\pm_{r}\Delta(r)\sqrt{\mathcal{R}(r)}}\mathrm{d}r$ (30)
$I_{t}$ $=f_{r_{s}}^{r_{o}}\frac{r^{2}\Delta(r)+2Mre^{-k/r}\left(r^{2}+a^{2}-a\lambda\right)}{\pm_{r}\Delta(r)\sqrt{\mathcal{R}(r)}}\mathrm{d}r$ (31)
@@
$G_{\theta}$ $=f_{\theta_{s}}^{\theta_{o}}\frac{\mathrm{~{}d}\theta}{\pm_{\theta}\sqrt{\Theta(\theta)}}$ (32)
$G_{\phi}$ $=f_{\theta_{s}}^{\theta_{o}}\frac{\csc^{2}\theta}{\pm_{\theta}\sqrt{\Theta(\theta)}}\mathrm{d}\theta$ (33)
$G_{t}$ $=f_{\theta_{s}}^{\theta_{o}}\frac{\cos^{2}\theta}{\pm_{\theta}\sqrt{\Theta(\theta)}}\mathrm{d}\theta$ (34)
where the integral notation $f$ means the path integrals along the photon trajectory, and these path integrals are monotonically increasing along the trajectory. We could see that the angular integrals of rotating regular black hole are same to Kerr black hole. However, the radial integrals are significantly different. For Kerr black hole, the radial integrals are analytical and revertible, however, we can only computer numerically for the radial integrals of rotating regular black hole due to its complicity. By inverting these integrals, we are able to know the source coordinates given a photon with conservative quantities $(\lambda,\eta)$.
Once we know the source coordinates, we can obtain the corresponding flux at this spacetime point according to the flux profile, and then adding photons accordingly to the image plane at position $(\alpha,\beta)$. Finally, we could obtain the black hole bolometric image. In this work, we consider a stationary axisymmetric thin accretion disk, so the emitting source only depends on the radius $r_{s}$. Every source ring with $r_{s}$ at the equatorial source plane (or the thin accretion disk) could have multiple images in the image plane, since they could be lensed multiple times by the rotating regular black hole. If a photon has cross the equatorial plane with $n$ times along its trajectory between the source and observer, then it will arrive at the image plane within a specific region where we called $n$th lensing band. For example, $n=0$ lensing band is just the direct image without any equatorial crossing in their trajectory. For the photons which have equatorial crossing once, it will located into the $n=1$ lensing band. The large $n$, the dimmer and narrower of the lensing bands, the closer to the critical curve [36, 37]. Therefore, we need to add up the photons in all the lensing bands to get a composite images of the black hole. We will only consider the lensing bands up to $n=2$ since it is almost coincide with the critical curve and much dimmer than the leading lensing bands. According to the radiative flux at the source coordinates, we can obtain the observed flux in the image plane according to the following relation,
$\mathcal{F}_{0}(\alpha,\beta)=\sum_{n=0}^{2}\chi^{4}\left(r_{\mathrm{s}}^{(n)},\alpha,\beta\right)\mathcal{F}_{\mathrm{s}}\left(r_{\mathrm{s}}^{(n)}\right)$ (35)
@@
$\chi=\frac{1}{u^{t}(1-\lambda\Omega)}$ (36)
@@ IV.2 Bolometric images of rotating regular black hole
By employing the above ray-tracing method, we modified the well-known aart [^1] https://github.com/iAART/aart code [53] allowing us to compute the observables numerically, and thus extend the code applicability to non-Kerr black holes. We investigate the lensing bands and bolometric images of rotating regular black holes as functions of black hole spin, the regular parameter, and the inclination angle. Furthermore, we perform a comparative analysis the results of rotating regular black hole with Kerr black holes, highlighting the differences induced by the presence of the regular parameter.
@@
Figure 6: Comparison of black hole images for Kerr black holes (left column) and regular black holes with $k=0.1$ (right column). The first to fourth rows correspond to parameters $(a=0.2,i=20^{\circ}),(a=0.8,i=20^{\circ}),(a=0.2,i=80^{\circ}),\text{and}(a=0.8,i=80^{\circ})$, respectively. The color bars in each subplot represent the brightness of the black hole image, with brighter regions corresponding to higher intensity values. The source flux profile $\mathcal{F}_{s}(r)$ for generating the observed image is taken from Fig.2 by making $\dot{m}=1$ and multiplying the values with a factor of $10^{5}$.
From Fig. 6, we could gain an intuitive understanding on the difference between the Kerr black hole and rotating regular black hole. As we discussed earlier, the shrinking effects in the rotating regular black hole are evident. Consequently, the structure of the rotating regular black hole becomes more concentrated compared to the Kerr black hole. Additionally, the rotating regular black holes display steeper and brighter brightness distributions than the Kerr black hole, indicating the influence of the modified geometry with $k=0.1$.
As shown in Fig. 2, the features of the thin accretion disk are also evident for different black hole spins and the regular parameter. Specifically, the inner edge of the accretion disk ($r_{ISCO}$) is smaller, and the radiative flux values are higher for larger black hole spins and regular parameters.
@@ V Conclusion and discussion
Rotating regular black holes, as a valuable exploration beyond GR, provide an intriguing model by resolving spacetime singularities. To test the physical implications of this idea, we investigated the observational features of thin accretion disks around this rotating regular black holes. Specifically, we analyzed key quantities of both time-like and null-like geodesics, illustrating how the regular parameter affects black hole geometrical properties such as horizons, the photon shell, and the ISCO. Based on these foundational results, we further examined the flux, temperature, and differential luminosity of the accretion disk as functions of the regular parameter and black hole spin. Additionally, using ray-tracing techniques, we explored the lensing behavior of the black hole and constructed realistic images of its surroundings, which shows a significant deviation from the standard Kerr metric in GR by comparing their images.

### v2

@@ IV.1 Ray-tracing method
In order to obtain the black hole image, we will apply the ray-tracing method. By shooting light rays from an observer’s image plane backward to the vicinity of the black hole, we could compute the resulting intensity in the image plane based on how the light rays intersect with the emission source (or the thin accretion disk) around the black holes.
@@
$\alpha=-\frac{\lambda}{\sin\theta_{o}},\quad\beta=\pm_{o}\sqrt{\eta+a^{2}\cos^{2}\theta_{o}-\lambda^{2}\cot^{2}\theta_{o}},$ (24)
@@
$\lambda=-\alpha\sin\theta_{o},\quad\eta=\left(\alpha^{2}-a^{2}\right)\cos^{2}\theta_{0}+\beta^{2},$ (25)
The conserved quantities corresponding to photon shells or bounded orbits $(\tilde{\lambda},\tilde{\eta})$ (see Eq. 17-19) will result in the critical curve $\mathcal{C}=\{\tilde{\alpha},\tilde{\beta}\}$ in the image plane since they are not stable. The enclose region by this critical curve in the image plane is usually denoted as the Black Hole Shadow. This critical curve will branch the behavior of light rays when we shoot backward from the image plane to the rotating regular black hole. For the light rays inside the critical curve, they will fall into black hole. While the light rays outside the critical curve are all scattered back to infinity, which means that they have a turning point in their trajectory. The radial turning point is the largest root of the radial potential $r_{4}$, and it is greater than the event horizon.
The relation Eq. (25) is not enough to determine the source coordinates of a photon. For a given black hole spin $a$, observer radial location $r_{o}$ and inclination angle $\theta_{o}$, our goal is to know the relation between the coordinates of the image plane $(\alpha,\beta)$ and the source coordinates $(r_{s},\theta_{s},\phi_{s},t_{s})$ where this photon emitted from. To do so, we need parameterize the photon trajectories by integrating the geodesic equations (3-6). In reference to [38, 39], the differential geodesic equation in rotating regular black hole spacetime could also recast into an integral form,
@@
$\Delta\phi\equiv\phi_{o}-\phi_{s}$ $=I_{\phi}+\lambda G_{\phi},$ (27)
$\Delta t\equiv t_{o}-t_{s}$ $=I_{t}+a^{2}G_{t},$ (28)
@@
$I_{r}$ $=f_{r_{s}}^{r_{o}}\frac{\mathrm{~{}d}r}{\pm_{r}\sqrt{\mathcal{R}(r)}},$ (29)
$I_{\phi}$ $=f_{r_{s}}^{r_{o}}\frac{a(2Mre^{-k/r}-a\lambda)}{\pm_{r}\Delta(r)\sqrt{\mathcal{R}(r)}}\mathrm{d}r,$ (30)
$I_{t}$ $=f_{r_{s}}^{r_{o}}\frac{r^{2}\Delta(r)+2Mre^{-k/r}\left(r^{2}+a^{2}-a\lambda\right)}{\pm_{r}\Delta(r)\sqrt{\mathcal{R}(r)}}\mathrm{d}r,$ (31)
@@
$G_{\theta}$ $=f_{\theta_{s}}^{\theta_{o}}\frac{\mathrm{~{}d}\theta}{\pm_{\theta}\sqrt{\Theta(\theta)}},$ (32)
$G_{\phi}$ $=f_{\theta_{s}}^{\theta_{o}}\frac{\csc^{2}\theta}{\pm_{\theta}\sqrt{\Theta(\theta)}}\mathrm{d}\theta,$ (33)
$G_{t}$ $=f_{\theta_{s}}^{\theta_{o}}\frac{\cos^{2}\theta}{\pm_{\theta}\sqrt{\Theta(\theta)}}\mathrm{d}\theta,$ (34)
where the integral notation $f$ means the path integrals along the photon trajectory, and these path integrals are monotonically increasing along the trajectory. We could see that the angular integrals of rotating regular black hole are same to Kerr black hole [38, 39]. However, the radial integrals are significantly different. For Kerr black hole, the radial integrals are analytical and revertible. However, it is not the case for rotating regular black hole, and we can only compute these radial integrals numerically. Anyway, by inverting these integrals, we are able to know the source coordinates given a photon with conservative quantities $(\lambda,\eta)$.
Once we know the source coordinates, we can obtain the corresponding flux at this spacetime point according to the emission profile, and then adding photons accordingly to the image plane at position $(\alpha,\beta)$. Finally, we could obtain the black hole bolometric image. In this work, we consider a stationary axisymmetric thin accretion disk, so the emitting source only depends on the radius $r_{s}$. Every source ring with $r_{s}$ at the equatorial source plane (or the thin accretion disk) could have multiple images in the image plane, since they could be lensed multiple times by the rotating regular black hole. If a photon has cross the equatorial plane with $n$ times along its trajectory between the source and observer, then it will arrive at the image plane within a specific region where we called $n$th lensing band. For example, $n=0$ lensing band is just the direct image without any equatorial crossing in their trajectory. For the photons which have equatorial crossing once, it will located into the $n=1$ lensing band. The large $n$, the dimmer and narrower of the lensing bands, the closer to the critical curve [38, 39]. Therefore, we need to add up the photons in all the lensing bands to get a composite images of the black hole. We will only consider the lensing bands up to $n=2$ since it is almost coincide with the critical curve and much dimmer than the leading lensing bands. According to the radiative flux profile at the source coordinates, we can obtain the observed flux in the image plane according to the following relation,
$\mathcal{F}_{0}(\alpha,\beta)=\sum_{n=0}^{2}\chi^{4}\left(r_{\mathrm{s}}^{(n)},\alpha,\beta\right)\mathcal{F}_{\mathrm{s}}\left(r_{\mathrm{s}}^{(n)}\right),$ (35)
@@
$\chi=\frac{1}{u^{t}(1-\lambda\Omega)},$ (36)
@@ IV.2 Bolometric images of rotating regular black hole
By employing the above ray-tracing method, we modified the well-known aart [^1] https://github.com/iAART/aart code [59] to allow us to compute the observables numerically, and thus extend the code applicability to non-Kerr black holes. We investigate the lensing bands and bolometric images of rotating regular black holes as functions of black hole spin, the regular parameter, and the inclination angle. Furthermore, we perform a comparative analysis the results of rotating regular black hole with Kerr black hole, highlighting the differences induced by the presence of the regular parameter.
@@
Figure 6: Comparison of black hole images for Kerr black holes (left column) and regular black holes with $k=0.1$ (right column). The first to fourth rows correspond to parameters $(a=0.2,i=20^{\circ}),(a=0.8,i=20^{\circ}),(a=0.2,i=80^{\circ})$,and $(a=0.8,i=80^{\circ})$, respectively. The color bars in each subplot represent the brightness of the black hole image, with brighter regions corresponding to higher bolometric values. The source flux profile $\mathcal{F}_{s}(r)$ for generating the observed images are taken from Fig. 2 by making $\dot{m}=1$ and multiplying the values with a factor of $10^{5}$.
From Fig. 6, we could gain an intuitive understanding on the differences between the Kerr black hole and rotating regular black hole. As we discussed earlier, the shrinking effects in the rotating regular black hole are evident. Consequently, the structure of the rotating regular black hole becomes more concentrated compared to the Kerr black hole. Additionally, the rotating regular black holes display steeper and brighter brightness distributions than the Kerr black hole, indicating the influence of the modified geometry with $k=0.1$.
As already shown in Fig. 2, the features of the thin accretion disk are also evident for different black hole spins and the regular parameter. Specifically, the inner edge of the accretion disk ($r_{ISCO}$) is smaller, and the radiative flux values are higher for larger black hole spins and regular parameters.
@@ V Conclusion and discussion
Rotating regular black holes, as a valuable exploration beyond GR, provide an intriguing model by resolving spacetime singularities. To test the physical implications of this idea, we investigated the observational features of thin accretion disks around this rotating regular black holes. Specifically, we analyzed key quantities of both time-like and null-like geodesics, illustrating how the regular parameter affects black hole geometrical properties such as horizons, the photon shell, and the ISCO. Based on these foundational results, we further examined the flux, temperature, and differential luminosity of the thin accretion disk as functions of the regular parameter and black hole spin. Additionally, using ray-tracing techniques, we explored the lensing behavior of the black hole and constructed realistic images of its thin accretion disk, which shows a significant deviation from the standard Kerr metric in GR by comparing their images.